# Retrieval-Enhanced Contrastive Vision-Text Models

### v1

@@ 4.2 Zero-shot transfer
Image classification and image/text retrieval. In Tab. 1, we observe that RECO allows to boost the zero-shot performance of CLIP on zero-shot image classification and retrieval, with large improvements especially on the fine-grained datasets. For example, we improve the original CLIP-B/32 accuracy by $+10.9$ on Cars, $+10.2$ on CUB and $+5.8$ on Flowers. The performance is also improved on less fine-grained benchmarks such as ImageNet or Places, though by more moderate margins (i.e. respectively $+1.1$ and $+1.6$). Secondly, we see in Tab. 1 that the performance gains are consistent across the two considered vision-text backbones (CLIP-B/32 and CLIP-L/14). For reference, we also report in Tab. 1 the numbers from other popular vision-text approaches [55, 27, 66, 8]. Note that our work is orthogonal to the choice of vision-text model and we expect that combining it with other powerful backbones, e.g. ALIGN [27], OpenCLIP [9] or CoCa [65], has the potential to bring further performance gains, especially in the fine-grained settings where their performance might be sub-optimal. Overall, the experiment in Tab. 1 confirms our initial motivation that retrieval from an external memory improves zero-shot recognition tasks, especially in fine-grained settings.
Open-domain visual entity recognition (OVEN). In Tab. 2, we show the zero-shot performance of RECO on the OVEN benchmark. We see that our method improves greatly over CLIP-L/14 on this challenging task, with an impressive relative improvement of $+138\%$. Note that we do not train or fine-tune our model on the OVEN training set. Remarkably, we observe in Tab. 2 that RECO also significantly outperforms much bigger models which are directly fine-tuned for this task, for example CLIP2CLIP [23] or PaLI-3B [8] while using respectively 2 $\times$ and 7 $\times$ less parameters. It even comes close to the performance of PaLI-17B while being 39 $\times$ smaller and not using any fine-tuning.
@@ 4.3 Design choice analyses
Uni-modal search and cross-modal fusion. In Tab. 3, we evaluate different alternatives for our method, namely (i) performing cross-modal search in the memory instead of uni-modal search and (ii) fusing uni-modal items (i.e. combining text with text and image with image) instead of cross-modal fusion. These different scenarios (uni- versus cross- modal search and fusion) are detailed in Section 3.1 and conceptually illustrated in Fig. 2. Firstly, we observe in Tab. 3 that uni-modal search (row 1) leads to a better performance compared to cross-modal search (row 2), with $+9.0$ versus $+1.7$ average relative improvement over CLIP. We remark that the gap is especially important for fine-grained datasets such as Cars, CUB and Flowers. This agrees with our hypothesis that cross-modal search suffers from the pre-trained CLIP model’s lack of fine-grained alignment between different modalities. By contrast, using the inherent strength of image and text representations within their respective modalities allows to retrieve relevant matches, as qualitatively observed in Fig. 4.
Secondly, we observe in Tab. 3 that uni-modal fusion (rows 3 and 4) works substantially worse than cross-modal fusion (rows 1 and 2). Indeed, we see that augmenting text embeddings with other text embeddings and image embeddings with other image embeddings does not bring any significant improvement over the baseline, and even tends to hurt the performance. Intuitively, a possible explanation is that cross-modal fusion allows us to inject complementary signal into the original embeddings [25]. By contrast, uni-modal provides signal that is already similar to the input, hence not as much additional information. Finally, we see in Tab. 3 that all the variants (rows 5, 6, 7 and 8) fail when simply averaging retrieved and original embeddings instead of learning the fusion with a transformer. This highlights the importance of learning to incorporate the retrieved items to the original embeddings before deploying the model at inference.
@@
Is the performance boost merely due to additional training? We validate the hypothesis that the performance gains are indeed due to our method, and not due to training an additional layer on top of CLIP. We replace RECO with an MLP layer of the same capacity initialized from scratch. We train it in a CLIP-style manner on the subset of Webli that we use when training RECO. We denote this subset by $\mathcal{R}_{\text{Webli}}$: it contains the $k=10$ nearest-neighbors for each CC12M datapoint retrieved from the Webli dataset, and contains 61M unique examples. Results are presented in Fig. 3 (left). We observe that training an extra layer on top of CLIP does not bring any performance gains and even deteriorates its performance. This is somehow expected since CLIP was extensively trained on a large and rich dataset [50] and additional training on a relatively small dataset deteriorates the general-purpose property of its representations. Overall, this experiment validates our retrieval-augmented approach.
@@ 5 Conclusion
In this paper, we introduce RECO, a method that enhances the fine-grained recognition capabilities of pre-trained vision-text models. Our approach shows the importance of uni-modal retrieval, yet cross-modal fusion for image and text inputs. We show that RECO consistently improves the performance on 11 different zero-shot tasks and that the gains are especially important in challenging fine-grained tasks. We include a broader impact discussion in Appendix.
Limitations. A limitation of this work is that it assumes to have access to a large and rich source of image-text pairs knowledge. While we show in Appendix that public datasets , e.g. LAION [53], can serve this purpose, the best of performance is obtained with a large private memory. Alternatively, one could use search engine APIs as the memory. Another limitation is that the performance gains of RECO come at the cost of increased inference time. In practice, we use a highly-optimized approximate $k$-NN algorithm [20], and querying $1$B examples in Webli takes milliseconds. It is also possible to mitigate this issue by reducing the number of retrieved items. As shown in Fig. 3, retrieving a single element already improves over CLIP substantially, while reducing compute requirements.
A Additional results
A.1 Using LAION-400M as the memory
In Table 5, we show that our method also works when using a public dataset as the memory bank instead of our private source of knowledge. Indeed, we observe that using LAION-400M [53] as the memory bank for RECO gives substantial gains of performance compared to the CLIP baseline across our different zero-shot tasks: for example +6.9 on Cars and +6.1 on CUB. This validates that our method is generic and can work with different choices of external knowledge.
Table 5: Choice of memory bank. We report zero-shot top-1 accuracy on different image classification tasks. We evaluate RECO when using two different sources of knowledge: the non publicly available WebLI [8] dataset (our default) and the publicly available LAION-400M [53] dataset.
Memory bank Publicly available Cars CUB Flowers Im1k Pl365
@@ A.1 Using LAION-400M as the memory
WebLI [8] (default) 68.1 (+10.9) 63.0 (+10.2) 67.9 (+5.8) 64.6 (+1.1) 42.2 (+1.6)
LAION-400M [53] 64.1 (+6.9) 58.9 (+6.1) 63.7 (+1.6) 63.4 (-0.1) 42.3 (+1.7)
A.2 More complex fusion module
@@
A.3 End-to-end finetuning versus frozen backbone
@@ B.1 Evaluation datasets
We report the details of each image classification dataset that we use to evaluate our model. Note that we only use test or validation splits of each of these datasets, training sets are disregarded. Stanford Cars (“Cars”) [32] contains $8,041$ test images of $196$ fine-grained car classes. Each class name consists of make, model and year of a car, e.g. 2012 Tesla Model S or 2012 BMW M3 coupe. CUB-200-2011 (“CUB”) [59] consists of $5,794$ test images of $200$ bird species. The dataset contains additional annotations, such as part locations, binary attibutes, and bounding boxes, but we do not use any of them. Oxford Flowers (“Flowers”) [45] has $6,149$ test images of $102$ flower categories commonly found in the United Kingdom. The dataset images contain strong visual variations, such as scale, pose, and light changes. ImageNet-1k (“Im1k”) [52], or also referred to as ILSVRC 2012, contains $50,000$ validation images from $1,000$ classes. Class names are obtained from the synsets in WordNet [43], and come from a large variety of generic concepts. Places365 (“Pl365”) [68] has $36,500$ validation images from $365$ generic scene categories. Some examples of class names include living room, cottage, lecture room, pier etc. Finally, Stanford Dogs (“Dogs”) [31] consists of $8,580$ test images from $120$ dog breeds.
@@ B.2 Evaluation protocols
Zero-shot image classification. We follow the standard setup [50] of embedding each class name with the text encoder. We classify an image to the class which has the highest cosine similarity between the image embedding and the corresping class name embedding. We report top-1 accuracy in all the image classification benchmarks. There is no variance at zero-shot evaluation time since the inference for both text and vision encoders are deterministic.
@@
OVEN benchmark. OVEN benchmark [23] is created by combining 14 existing datasets (ImageNet21k-P [51, 52], iNaturalist2017 [57], Cars196 [32], SUN397 [64], Food101 [5], Sports100 [16], Aircraft [39], Oxford Flowers [45], Google Landmarks v2 [61], and various VQA (visual question answering) datasets [18, 69, 33, 40, 56, 16]) and grounding their categories to Wikipedia entities. The benchmark consists of two splits. Entity Split measures the image recognition or retrieval capabilities of a model, whereas the Query Split is designed as a VQA task. We focus on the Entity Split in this paper.
@@ C Illustrative comparison of uni-/cross- modal search and uni-/cross- fusion
D Broader Impact
We propose a retrieval-based recognition approach, where we search for similar images and text in a large-scale memory. Data retrieved from such uncurated sources may be biased against certain populations across the world [49, 10]. Furthermore, it is important that the privileged user data does not exist in such data collections, in order to avoid using the data without the consent of its owner. We acknowledge these potential misuses, and encourage the community to utilize more fair and responsible data collections.

### v2

@@ 4.2 Zero-shot transfer
Image classification. In Tab. 1, we observe that RECO boosts the zero-shot performance of CLIP and LiT on zero-shot image classification with large improvements especially on the fine-grained datasets. For example, we improve the original CLIP-B/32 accuracy by $+10.9$ on Cars, $+10.2$ on CUB and $+5.8$ on Flowers. The performance is also improved on less fine-grained benchmarks such as ImageNet or Places, though by more moderate margins (i.e. respectively $+1.1$ and $+1.6$). Secondly, we see in Tab. 1 that the performance gains are consistent across all vision-text backbones (CLIP-R-50, CLIP-B/32, CLIP-L/14, and LiT-L16L). Note that LiT-L16L is pre-trained on Webli, which is our memory bank, and we still observe the benefits of RECO. For reference, we also report in Tab. 1 the numbers from other popular vision-text approaches (Jia et al., 2021; Chen et al., 2023). Overall, the experiment in Tab. 1 confirms our initial motivation that retrieval from an external memory improves zero-shot recognition tasks, especially in fine-grained settings.
Open-domain visual entity recognition (OVEN). In Tab. 2, we show the zero-shot performance of RECO on the OVEN benchmark. We see that our method improves greatly over CLIP-L/14 on this challenging task, with an impressive relative improvement of $+132\%$. Note that we do not train or fine-tune our model on the OVEN training set. Remarkably, we observe in Tab. 2 that RECO also significantly outperforms much bigger models which are directly fine-tuned for this task, for example CLIP2CLIP (Hu et al., 2023) or PaLI-3B (Chen et al., 2023) while using respectively 2 $\times$ and 7 $\times$ less parameters. It even comes close to the performance of PaLI-17B while being 39 $\times$ smaller and not using any fine-tuning.
@@ 4.3 Design choice analyses
Uni-modal search and cross-modal fusion. In Tab. 3, we evaluate different alternatives for our method, namely (i) performing cross-modal search in the memory instead of uni-modal search and (ii) fusing uni-modal items (i.e. combining text with text and image with image) instead of cross-modal fusion. These different scenarios (uni- versus cross- modal search and fusion) are detailed in Section 3.1 and conceptually illustrated in Fig. 2. Firstly, we observe in Tab. 3 that uni-modal search (row 1) leads to a better performance compared to cross-modal search (row 2), with $+9.0$ versus $+1.7$ average relative improvement over CLIP. We remark that the gap is especially important for fine-grained datasets such as Cars, CUB and Flowers. This agrees with our hypothesis that cross-modal search suffers from the pre-trained CLIP model’s lack of fine-grained alignment between different modalities. By contrast, using the inherent strength of image and text representations within their respective modalities allows to retrieve relevant matches, as qualitatively observed in Fig. 4. Secondly, we observe in Tab. 3 that uni-modal fusion (rows 3 and 4) works substantially worse than cross-modal fusion (rows 1 and 2). Indeed, we see that augmenting text embeddings with other text embeddings and image embeddings with other image embeddings does not bring any significant improvement over the baseline, and even tends to hurt the performance. Intuitively, a possible explanation is that cross-modal fusion allows us to inject complementary signal into the original embeddings (Iscen et al., 2023). By contrast, uni-modal provides signal that is already similar to the input, hence not as much additional information. Finally, we see in Tab. 3 that all the variants (rows 5, 6, 7 and 8) fail when simply averaging retrieved and original embeddings instead of learning the fusion with a transformer. This highlights the importance of learning to incorporate the retrieved items to the original embeddings before deploying the model at inference.
@@
Is the performance boost merely due to additional training? We replace RECO with an MLP layer of the same capacity initialized from scratch. We train it in a CLIP-style manner on the subset of Webli that we use when training RECO. We denote this subset by $\mathcal{R}_{\text{Webli}}$: it contains the $k=10$ nearest-neighbors for each CC12M datapoint retrieved from the Webli dataset, and contains 61M examples. We observe in Fig. 3 (left) that training an extra layer on top of CLIP does not bring any gains and even deteriorates its performance. Indeed, CLIP was extensively trained on a large dataset (Radford et al., 2021) and additional training on a relatively small dataset deteriorates the general-purpose property of its representations. Overall, this experiment validates that the performance gains are due to our method and not to training an additional layer on top of CLIP.
@@
Limitations. A limitation of this work is that it assumes to have access to a large and rich source of image-text pairs knowledge. While we show in Appendix that public datasets , e.g. LAION (Schuhmann et al., 2021), can serve this purpose, the best of performance is obtained with a large private memory. Alternatively, one could use search engine APIs as the memory. Another limitation is that the performance gains of RECO come at the cost of increased inference time. In practice, we use a highly-optimized approximate $k$-NN algorithm (Guo et al., 2020). It takes about 14ms to query Webli (956M examples) with a single 512-d ViT-B/32 CLIP embedding. Using retrieval at inference time incurs an overhead of $25\%$ compared to not using any retrieval at inference time (e.g. baseline CLIP model), but improves the accuracy by up to 10.9.
@@ 5 Conclusion
In this paper, we introduce RECO, a method that enhances the fine-grained recognition capabilities of pre-trained vision-text models. Our approach shows the importance of uni-modal retrieval, yet cross-modal fusion for image and text inputs. We show that RECO consistently improves the performance on 11 zero-shot tasks and that the gains are especially important in challenging fine-grained tasks.
A Appendix
A.1 Image/text retrieval
Tab. 5 shows that RECO allows to boost the zero-shot performance of CLIP for image and text retrieval even further. We observe that we can improve the recall@1 up to $4.6$ points for text-to-image retrieval. We also see smaller, but consistent gains for image-to-text retrieval.
A.2 Using LAION-400M as the memory
In Table 6, we show that our method also works when using a public dataset as the memory bank instead of our private source of knowledge. Indeed, we observe that using LAION-400M (Schuhmann et al., 2021) as the memory bank for RECO gives substantial gains of performance compared to the CLIP baseline across our different zero-shot tasks: for example +6.9 on Cars and +6.1 on CUB. This validates that our method is generic and can work with different choices of external knowledge.
Table 5: Zero-shot transfer to retrieval. We report recall@1 for retrieval. We show the improvements obtained with RECO on top of CLIP-B/32 and CLIP-L/14: absolute performance gains are between brackets. For reference, we also include the performance of other standard image-text foundation models (Flava (Singh et al., 2022) and Align-base (Jia et al., 2021). We also report the total parameter count (“# par.”) of the different models (in Million).
[…]
Method # par. COCO Flickr COCO Flickr
CLIP-B/32 151 30.2 61.1 51.2 80.9
  + RECO 154 33.6(+3.4) 65.7(+4.6) 52.2(+1.1) 81.8(+0.9)
CLIP-L/14 428 35.2 68.6 57.2 87.5
  + RECO 435 38.7(+3.5) 72.6(+4.0) 58.0(+0.8) 88.5(+1.0)
Other approaches
Flava 172 38.4 65.2 42.7 67.7
Align 247 40.2 72.6 55.1 86.7
Table 6: Choice of memory bank. We report zero-shot top-1 accuracy on different image classification tasks. We evaluate RECO when using two different sources of knowledge: the non publicly available WebLI (Chen et al., 2023) dataset (our default) and the publicly available LAION-400M (Schuhmann et al., 2021) dataset.
Memory bank Public Cars CUB Flowers Im1k Pl365
@@ A.2 Using LAION-400M as the memory
WebLI (Chen et al., 2023) (default) 68.1 (+10.9) 63.0 (+10.2) 67.9 (+5.8) 64.6 (+1.1) 42.2 (+1.6)
LAION-400M (Schuhmann et al., 2021) 64.1 (+6.9) 58.9 (+6.1) 63.7 (+1.6) 63.4 (-0.1) 42.3 (+1.7)
A.3 More complex fusion module
@@
A.4 End-to-end finetuning versus frozen backbone
@@ B.1 Evaluation datasets
We report the details of each image classification dataset that we use to evaluate our model. Note that we only use test or validation splits of each of these datasets, training sets are disregarded. Stanford Cars (“Cars”) (Krause et al., 2013) contains $8,041$ test images of $196$ fine-grained car classes. Each class name consists of make, model and year of a car, e.g. 2012 Tesla Model S or 2012 BMW M3 coupe. CUB-200-2011 (“CUB”) (Wah et al., 2011) consists of $5,794$ test images of $200$ bird species. The dataset contains additional annotations, such as part locations, binary attibutes, and bounding boxes, but we do not use any of them. Oxford Flowers (“Flowers”) (Nilsback & Zisserman, 2008) has $6,149$ test images of $102$ flower categories commonly found in the United Kingdom. The dataset images contain strong visual variations, such as scale, pose, and light changes. ImageNet-1k (“Im1k”) (Russakovsky et al., 2015), or also referred to as ILSVRC 2012, contains $50,000$ validation images from $1,000$ classes. Class names are obtained from the synsets in WordNet (Miller, 1998), and come from a large variety of generic concepts. Places365 (“Pl365”) (Zhou et al., 2017) has $36,500$ validation images from $365$ generic scene categories. Some examples of class names include living room, cottage, lecture room, pier etc. Finally, Stanford Dogs (“Dogs”) (Khosla et al., 2011) consists of $8,580$ test images from $120$ dog breeds.
@@ B.2 Evaluation protocols
Zero-shot image classification. We follow the standard setup (Radford et al., 2021) of embedding each class name with the text encoder. We classify an image to the class which has the highest cosine similarity between the image embedding and the corresping class name embedding. We report top-1 accuracy in all the image classification benchmarks. There is no variance at zero-shot evaluation time since the inference for both text and vision encoders are deterministic.
@@
OVEN benchmark. OVEN benchmark (Hu et al., 2023) is created by combining 14 existing datasets (ImageNet21k-P (Ridnik et al., 2021; Russakovsky et al., 2015), iNaturalist2017 (Van Horn et al., 2018), Cars196 (Krause et al., 2013), SUN397 (Xiao et al., 2010), Food101 (Bossard et al., 2014), Sports100 (Gerry, 2021), Aircraft (Maji et al., 2013), Oxford Flowers (Nilsback & Zisserman, 2008), Google Landmarks v2 (Weyand et al., 2020), and various VQA (visual question answering) datasets (Goyal et al., 2017; Zhu et al., 2016; Krishna et al., 2017; Marino et al., 2019; Singh et al., 2019; Gerry, 2021)) and grounding their categories to Wikipedia entities. The benchmark consists of two splits. Entity Split measures the image recognition or retrieval capabilities of a model, whereas the Query Split is designed as a VQA task. We focus on the Entity Split in this paper.
@@ C Illustrative comparison of uni-/cross- modal search and uni-/cross- fusion
D Near-Duplicate Filtering
Figure 8: Left: Cosine similarity distribution between the test image queries of CUB2011 and their 10 kNN. Right: Some of the queries and their nearest neighbors from the memory with $0.95$ similarity.
Figure 8 (Left) of the PDF shows the cosine similarity distribution between the queries and memory images on CUB2011 dataset. It is shown that most of the retrieved items have a similarity around 0.9, which indicates that most of the retrieved elements are similar but not identical. Note that there is no cosine similarity above 0.95. This is because we remove any memory image (and their corresponding caption) if they have a similarity of 0.95 or above with any of the test images. Figure 8 (Right) shows some of the kNNs with 0.95 cosine similarity. We see that the retrieved examples are not near-duplicates, but they are conceptually similar as they should be.
E Broader Impact
We propose a retrieval-based recognition approach, where we search for similar images and text in a large-scale memory. Data retrieved from such uncurated sources may be biased against certain populations across the world (Prabhu & Birhane, 2020; De Vries et al., 2019). Furthermore, it is important that the privileged user data does not exist in such data collections, in order to avoid using the data without the consent of its owner. We acknowledge these potential misuses, and encourage the community to utilize more fair and responsible data collections.
F Noisy examples
We show the 10 retrieved captions for some of the test queries on the CUB2011 dataset on Figure 9. We observe that RECO predicts the correct class even if the retrieved captions contain irrelevant examples. This is because the Transformer module learns to aggregate the retrieved captions, implicitly keeping the relevant ones and disregarding the irrelevant ones. Note that Webli contains captions in languages other than English, which are considered noisy for CLIP, because it is trained in an English corpus.
Figure 9: Noisy examples. We display image queries and their labels on the left panel and 10 retrieved captions and images on the right panel. Retrieved captions are in green if they contain useful information, red otherwise. RECO learns